OpenStudy (anonymous):

Simplify the following expression -sqrt32/49

OpenStudy (jack1):

-sqrt 32/49 = -sqrt (16) * sqrt (2) / 49 = -4 sqrt (2) / 49 = -4 sqrt (2) / 7^2